In a group of people, 27% are male. In the same group, 41% commute to work by using public transportation. What percentage of females in this group commute by some other form of transportation (other than public)? It is safe to assume that form of transportation is independent of sex. Round all calculated probabilities (including final answer) to two decimal places throughout your solution process.
Answer #1

If 27% people in a group are male, so 100-27 = 73% are female.

41% people in the group commute by public transportation, so 100-41= 59% people commute by other forms of transportation.

Now, out of 73% females, 59% commute by other form of transportation, so required percentage is 73*59 = 43.07%

In Calculation of the above probability, we are multiplying the percentage of females in the group with the percentage of people who travel by other forms of transportation (which is the basic rules of Independence of two events), so yes we have assumed that transportation is independent of sex here.